# Exponentially accelerated relaxation and quantum Mpemba effect in open quantum systems

### v1

@@ Abstract
We investigate the quantum Mpemba effect in the relaxation of open quantum systems whose effective dynamics is described by Davies maps. We present a class of unitary transformations based on permutation matrices that, acting on the initial state of the system, (i) suppress the contribution of slowest decaying modes of the nonunitary dynamics; (ii) ensure that it is as distinguishable as possible from the steady state. The first requirement guarantees an exponentially accelerating convergence to the steady state, while the second implies that a quantum system initially farther from equilibrium approaches it more rapidly than an initial state closer to it. This protocol provides a genuine Mpemba effect, and its numerical simulation requires low computational effort. We prove that, for any initial state, there always exists a permutation matrix that maximizes its distance from the equilibrium for a given information-theoretic distinguishability measure. We illustrate our findings for the nonunitary dynamics of the transverse field Ising chain and XXZ chain, each weakly coupled to a bosonic thermal bath, showing the quantum Mpemba effect captured by the Hilbert-Schmidt distance, quantum relative entropy, and trace distance. Our results provide a universal and versatile framework to engineer the genuine quantum Mpemba effect in open quantum systems.
@@ I Introduction
Intuition suggests that the cooling of a physical system is a monotonic process: the closer to the equilibrium, the faster it should reach it. However, observations dating back to Aristotle over 2000 years ago Aristotle (1952), as well as studies developed by Mpemba and Osborne Mpemba and Osborne (1969), and Kell Kell (1969) in 1969, reveal counterintuitive behavior: under certain conditions, initially hotter systems can cool more rapidly than initially colder systems. This effect is called as the Mpemba effect. Since then, the phenomenon has been reported in different physical contexts, including crystalline polymers, manganites, and colloidal suspensions Hu et al. (2018); Chaddah et al. ; Kumar and Bechhoefer (2020). Despite its widespread occurrence, the underlying mechanism remains the subject of intense debate Burridge and Linden (2016); Bechhoefer et al. (2021). Recent advances in stochastic thermodynamics suggest that initially hotter systems can exploit dynamical shortcuts, accessing regions of state space that allow them to relax more rapidly than colder systems Lu and Raz (2017); Klich et al. (2019). We also mention the interplay between Mpemba and Kovacs effects in the so-called time-delayed cooling law Santos (2024).
@@
The Mpemba effect has found an analogue in the quantum regime Nava and Fabrizio (2019). The so-called quantum Mpemba effect (QME) is defined as a physical process where a system initially further from equilibrium relaxes faster than a system closer to equilibrium. In Markovian open quantum systems, by acting a suitable unitary operator on the pure initial state of the system, one verifies the suppression of the slowest decay mode of the nonunitary dynamics Carollo et al. (2021); Kochsiek et al. (2022, 2022); Bao and Hou (2025). This mechanism ensures an exponentially faster convergence to the stationary state, which defines the QME. It is worth noting, however, that the slowest decaying mode does not always contribute to the relaxation mechanism. This was observed, for example, in the effective Markovian dynamics of a single-level quantum dot coupled to two reservoirs Chatterjee et al. (2023). We refer to Refs. Ares et al. (2025a); Teza et al. for recent reviews on QME.
@@
The quantum Mpemba effect also manifests itself in the restoration of broken symmetries in many-body quantum systems. In this setting, QME occurs when the symmetry is locally restored faster for the initial state the breaks it more Rylands et al. (2024); Ares et al. (2023); Liu et al. (2024); Yamashika et al. (2024); Ares et al. (2025b); Yamashika et al. (2025); Russotto et al. (2025). We also mention recent works addressing QME in the study of quantum complexity Aditya et al. ; Mackinnon and Paternostro , local relaxation of closed quantum systems Ares et al. (2025c), and quantum speed limits Bettmann and Goold (2025). Experimental evidence of the quantum Mpemba effect was reported in trapped-ion quantum simulator Joshi et al. (2024); Zhang et al. (2025), including its inverse version Aharony Shapira et al. (2024), and more recently QME was addressed in a Nuclear Magnetic Resonance platform Schnepper et al. .
@@
In this work, we present a framework that allows for exponentially faster relaxation to equilibrium and QME in open quantum systems described by Davies maps [see Fig. 1(a)]. Our approach takes into account a unitary operator based on permutation matrices which, once applied to the initial state of the system, rearranges its spectrum and induces the suppression of the slowest decaying mode of the dynamics [see Fig. 1(b)]. On the one hand, we realize that exponential speed up is achieved regardless the permutation matrix. On the other hand, we prove that there always exists a permutation matrix that maximizes the distance from the initial state to equilibrium. This distance is characterized by paradigmatic distinguishability measures, namely, Hilbert-Schmidt distance, quantum relative entropy, and trace distance. Together, these features guarantee the occurrence of the genuine quantum Mpemba effect. We present numerical simulations to support our findings. We emphasize that our approach can also be appropriately modified to encompass other class of nonunitary evolutions, thus recovering previous results from the literature. Finally, we show that, in order to achieve an exponentially rapid evolution towards equilibrium, an initial state does not necessarily need to be incoherent respective to the eigenbasis of the Hamiltonian.
@@
The outline of the paper is as follows. In Sec. II, we revise the general theory of Davies maps, particularly addressing the spectral properties of the Liouvillian that governs the effective nounitary dynamics of the quantum system. In Sec. III, we present a framework that engineers an exponential speed up relaxation by eliminating the slowest decay modes of the Liouvillian. In Sec. IV, we show that there is always possible to find a permutation matrix such that we have a dressed state far from the initial state. Hence, once our protocol induces quicker relaxation towards equilibrium, these two features trigger genuine quantum Mpemba effect. In Sec. V, we illustrate our findings for the transverse field Ising model, and the XXZ model, each of these spin chains weakly coupled to a thermal bath. In Sec. VI, we showed that there are initial states for which an exponentially rapid relaxation to the equilibrium is already observed, but which are not incoherent with respect to the energy eigenbasis. Finally, in Sec. VII we summarize our conclusions.
@@ III Speeding up the relaxation process
The unitary transformation in Eq. (III) first maps $\rho({t_{0}})$ to its diagonal form. The permutation matrix $P_{\pi}$} then rearranges the order of the diagonal entries of $D$, which in turn is mapped onto the diagonal matrix ${P_{\pi}}{D}{P_{\pi}^{\dagger}}=\text{diag}({\lambda_{\pi(1)}},\ldots,{\lambda_{\pi(d)}})$. The permutation matrix $P_{\pi}$} is constructed in order to maximize the distinguishability between ${\rho^{\prime}}({t_{0}})$ and the steady state $\texttt{R}_{1}$. This means that the diagonal entries of $D$ will be rearranged in ascending order, ${\lambda_{\pi(1)}}<\ldots<{\lambda_{\pi(d)}}$, or even descending order, ${\lambda_{\pi(1)}}>\ldots>{\lambda_{\pi(d)}}$, depending on the figure of merit used to evaluate the distance from the transformed state to equilibrium. It will become clear later that $P_{\pi}$} plays an essencial role in ensuring that the genuine the Mpemba effect occurs for a given nonunitary dynamics [see Sec. IV]. Finally, the resulting matrix is rotated by $U_{1}$ to the quantum state in Eq. (III), which is incoherent with respect to the eigenbasis of the Hamiltonian $H$. In comparison with the results from Ref. Moroder et al. (2024), we note that the numerical simulation of our protocol requires low computational effort. The calculations involve prior knowledge of the spectrum of operators $H$ and $\rho({t_{0}})$, while the permutation matrix is determined by the appropriate ordering of the eigenvalues of $\rho({t_{0}})$ to maximize the distance from ${\rho^{\prime}}({t_{0}})$ to equilibrium.
@@
We observe that the operators $\mathcal{L}_{\text{vec}}$} and $\Xi$ share the same spectrum of $d^{2}$} eigenvalues, of which $d$ are real and come from the subblock $\Xi_{\text{off}}$}, and the remaining d(d1)d(d-1) eigenvalues are complex and related to $\Xi_{\text{diag}}$}. On the one hand, we find that the subblock $\Xi_{\text{off}}$} contributes with $d$ eigenvectors $|{\texttt{L}^{\prime}_{s}}\rangle\rangle$ and $|{\texttt{R}^{\prime}_{s}}\rangle\rangle$, whose matrix forms are Hermitian, fully diagonal eigenmatrices $\texttt{L}^{\prime}_{s}$} and $\texttt{R}^{\prime}_{s}$} respective to the basis $\{|{k}\rangle\}_{k=1,\ldots,d}$. This naturally includes the eigenoperator $\texttt{R}^{\prime}_{1}$} related to the steady state ${\texttt{R}_{1}}={U_{1}}{\texttt{R}^{\prime}_{1}}{U_{1}^{\dagger}}$ of the dynamics, and also the vector $|{\texttt{L}^{\prime}_{1}}\rangle\rangle$ constrained to the eigenmatrix $\texttt{L}_{1}$} that plays the role of the identity. On the other hand, the slowest decaying modes of the nonunitary dynamics are related to the subspace spanned by d(d1)d(d-1) eigenvectors of $\Xi_{\text{diag}}$}. Noteworthy, the respective set of left $|{\texttt{L}^{\prime}_{s}}\rangle\rangle$ and right $|{\texttt{R}^{\prime}_{s}}\rangle\rangle$ eigenvectors gives rise to upper/lower triangular matrices $\texttt{L}^{\prime}_{k}$ and $\texttt{R}^{\prime}_{s}$}, respectively, which have all entries equal to zero except for one off-diagonal element. In detail, one finds that $\langle{j}|{\texttt{L}^{\prime}_{s}}|{l}\rangle={\delta_{j,{j_{0}}}}{\delta_{l,{l_{0}}}}$, and $\langle{j}|{\texttt{R}^{\prime}_{s}}|{l}\rangle={\delta_{j,{j_{0}}}}{\delta_{l,{l_{0}}}}$, for a given pair ${j_{0}},{l_{0}}\in\{1,\ldots,d\}$, with ${j_{0}}\neq{l_{0}}$. In particular, we note that this set comprises the eigenoperator $\texttt{L}^{\prime}_{2}$} connected to the lowest decaying mode ${\texttt{L}_{2}}={U_{1}}{\texttt{L}^{\prime}_{2}}{U_{1}^{\dagger}}$ of the Davies map. It turns out that ${\texttt{L}^{\prime}_{2}}=|{j_{0}}\rangle\langle{l_{0}}|$ is a non-Hermitian, upper/lower triangular matrix.
@@ IV.1 Hilbert-Schmidt distance
${\mathcal{D}_{\text{HSD}}}(\rho,\varrho)=\sqrt{\text{Tr}(|\rho-\varrho|)}\penalty 10000\ ,$ (17)
with $|{X}|:=\sqrt{{X^{\dagger}}X}$. We note that HSD is (i) non-negative, i.e., ${\mathcal{D}_{\text{HSD}}}(\rho,\varrho)\geq 0$, with ${\mathcal{D}_{\text{HSD}}}(\rho,\varrho)=0$ if and only if $\rho=\varrho$; (ii) symmetric, ${\mathcal{D}_{\text{HSD}}}(\rho,\varrho)={\mathcal{D}_{\text{HSD}}}(\varrho,\rho)$; (iii) isometric invariant, ${\mathcal{D}_{\text{HSD}}}(V\rho{V^{\dagger}},V\varrho{V^{\dagger}})={\mathcal{D}_{\text{HSD}}}(\rho,\varrho)$, where ${V^{\dagger}}={V^{-1}}$ is a unitary operator. It is worth to note that HSD satisfy the triangle inequality, ${\mathcal{D}_{\text{HSD}}}(\rho,\varrho)\leq{\mathcal{D}_{\text{HSD}}}(\rho,\varpi)+{\mathcal{D}_{\text{HSD}}}(\varpi,\varrho)$, for all $\rho,\varpi,\varrho\in\mathcal{S}$. However, HSD it is not always contractive under completely positive and trace-preserving maps Ozawa (2000).
@@
where $A_{l}$ is a permutation matrix, $0\leq{\epsilon_{l}}\leq 1$, with ${\sum_{l}}\,{\epsilon_{l}}=1$, while $D$ and $\Sigma$ are diagonal matrices containing the spectrum of $\rho({t_{0}})$ and $\texttt{R}_{1}$}, respectively. The right-hand side of Eq. (19) defines a convex sum ${\sum_{l}}\,{\epsilon_{l}}{c_{l}}$ of the nonnegative elements ${c_{l}}:=\text{Tr}({A_{l}}{D}{A_{l}^{\dagger}}\Sigma)$. We note that ${c_{\text{min}}}\leq{\sum_{l}}\,{\epsilon_{l}}{c_{l}}\leq{c_{\text{max}}}$, where ${c_{\text{min}}}=\text{min}\{{c_{1}},{c_{2}},\ldots,{c_{d}}\}$ and ${c_{\text{max}}}=\text{max}\{{c_{1}},{c_{2}},\ldots,{c_{d}}\}$. This means that there exist an optimal permutation matrix $P_{\pi,\text{opt}}$} such that $\text{Tr}({P_{\pi,\text{opt}}}{D}{P_{\pi,\text{opt}}^{\dagger}}\Sigma)$ is closest to the minimum value $c_{\text{min}}$, i.e., $\text{Tr}({P_{\pi,\text{opt}}}{D}{P_{\pi,\text{opt}}^{\dagger}}\Sigma)\leq{c_{\text{min}}}$. To see this point, we note that
@@
where $\pi(\bullet)$ maps index $l$ to $\pi(l)$, with $l,\pi(l)\in\{1,2,\ldots,d\}$. Therefore, to ensure that ${\sum_{l=1}^{d}}\,{\alpha_{l}}{\lambda_{\pi(l)}}\leq{c_{\text{min}}}$, it suffices to choose a permutation matrix such that ${P_{\pi,\text{opt}}}{D}{P^{\dagger}_{\pi,\text{opt}}}=\text{diag}({\lambda_{\pi(1)}},{\lambda_{\pi(2)}},\ldots,{\lambda_{\pi(d)}})$, with ${\lambda_{\pi(1)}}\leq{\lambda_{\pi(2)}}\leq\ldots\leq{\lambda_{\pi(d)}}$ being listed in ascending order, provided that ${\alpha_{1}}\geq{\alpha_{2}}\geq\ldots\geq{\alpha_{n}}$ are listed in descending order, or vice-versa Hardy et al. (1988). This proves the validity of the inequality in Eq. (18), which means that there exists a permutation $P_{\pi}$ that maximizes HSD evaluated for states ${\rho^{\prime}}({t_{0}})$ and $\texttt{R}_{1}$}.

### v2

@@ Abstract
We investigate the quantum Mpemba effect in the relaxation of open quantum systems whose effective dynamics is described by Davies maps. We present a class of unitary transformations built from permutation matrices that, when applied to the initial state of the system, (i) suppress the slowest decaying modes of the nonunitary dynamics; (ii) maximize its distinguishability from the steady state. The first requirement guarantees exponentially accelerating convergence to the steady state, and the second implies that a quantum system initially farther from equilibrium approaches it more rapidly than one that starts closer. This protocol provides a genuine Mpemba effect, and its numerical simulation requires low computational effort. We prove that, for any initial state, one can always find a permutation matrix that maximizes its distance from equilibrium for a specified information-theoretic distinguishability measure. We illustrate our findings for a two-level system, and also for the nonunitary dynamics of the transverse field Ising chain and XXZ chain, each weakly coupled to a bosonic thermal bath, and demonstrate the quantum Mpemba effect as captured by the Hilbert-Schmidt distance, quantum relative entropy, and trace distance. Our results provide a versatile framework to engineer the genuine quantum Mpemba effect in Markovian open quantum systems.
@@ I Introduction
Intuition suggests that the cooling of a physical system is a monotonic process: the closer the system is to equilibrium, the faster it should reach that state. However, observations dating back to Aristotle over 2000 years ago Aristotle (1952), along with modern studies developed by Mpemba and Osborne Mpemba and Osborne (1969), and Kell Kell (1969) in 1969, reveal counterintuitive behavior: under certain conditions, initially hotter systems can cool more rapidly than initially colder systems. This phenomenon is known as the Mpemba effect. Since then, it has been observed in various physical systems, including crystalline polymers, manganites, and colloidal suspensions Hu et al. (2018); Chaddah et al. ; Kumar and Bechhoefer (2020). Despite its widespread occurrence, the underlying mechanism remains the subject of intense debate Burridge and Linden (2016); Bechhoefer et al. (2021). Recent advances in stochastic thermodynamics suggest that initially hotter systems can exploit dynamical shortcuts, accessing regions of state space that allow them to relax more rapidly than colder systems Lu and Raz (2017); Klich et al. (2019). We also mention the interplay between Mpemba and Kovacs effects in the so-called time-delayed cooling law Santos (2024).
@@
The Mpemba effect has found an analogue in the quantum regime Nava and Fabrizio (2019). The so-called quantum Mpemba effect (QME) is defined as a physical process where a system initially further from equilibrium relaxes faster than a system closer to equilibrium. In Markovian open quantum systems, by acting a suitable unitary operator on the pure initial state of the system, one verifies the suppression of the slowest decay mode of the nonunitary dynamics Carollo et al. (2021); Kochsiek et al. (2022, 2022); Bao and Hou (2025). This mechanism ensures an exponentially faster convergence to the stationary state, which defines the QME. It is worth noting, however, that the slowest decaying mode does not always contribute to the relaxation mechanism. This was observed, for example, in the effective Markovian dynamics of a single-level quantum dot coupled to two reservoirs Chatterjee et al. (2023). We refer to Refs. Ares et al. (2025a); Teza et al. ; Yu et al. (2025) for recent reviews on QME.
@@
The quantum Mpemba effect also manifests itself in the restoration of broken symmetries in many-body quantum systems. In this setting, QME occurs when the symmetry is locally restored faster for the initial state the breaks it more Rylands et al. (2024); Ares et al. (2023); Liu et al. (2024); Yamashika et al. (2024); Ares et al. (2025b); Yamashika et al. (2025); Russotto et al. (2025); Summer et al. . We also mention recent works addressing QME in the study of quantum complexity Aditya et al. ; Mackinnon and Paternostro , local relaxation of closed quantum systems Ares et al. (2025c), and quantum speed limits Bettmann and Goold (2025). Experimental evidence of the quantum Mpemba effect was reported in trapped-ion quantum simulator Joshi et al. (2024); Zhang et al. (2025), including its inverse version Aharony Shapira et al. (2024), and more recently QME was addressed in a Nuclear Magnetic Resonance platform Schnepper et al. .
@@
In this work, we present a framework that allows for exponentially faster relaxation to equilibrium and QME in open quantum systems described by Davies maps [see Fig. 1(a)]. Our approach takes into account a unitary operator based on permutation matrices which, once applied to the initial state of the system, rearranges its spectrum and induces the suppression of the slowest decaying mode of the dynamics [see Fig. 1(b)]. On the one hand, we realize that exponential speed up is achieved regardless the permutation matrix. On the other hand, we prove that there always exists a permutation matrix that maximizes the distance from the initial state to equilibrium. This distance is characterized by paradigmatic distinguishability measures, namely, Hilbert-Schmidt distance, quantum relative entropy, and trace distance. Together, these features guarantee the occurrence of the genuine quantum Mpemba effect. We present numerical simulations to support our findings. We emphasize that our approach can also be suitably adapted to encompass other classes of nonunitary evolutions, in particular recovering the main result of Ref. Carollo et al. (2021) [see Appendix A]. Finally, we show that achieving exponentially fast evolution toward equilibrium does not necessarily require the initial state to be incoherent with respect to the eigenstates of the Hamiltonian.
@@
The outline of the paper is as follows. In Sec. II, we revise the general theory of Davies maps, particularly addressing the spectral properties of the Liouvillian that governs the effective nounitary dynamics of the quantum system. In Sec. III, we present a framework that engineers an exponential speed up relaxation by eliminating the slowest decay modes of the Liouvillian. In Sec. IV, we show that there is always possible to find a permutation matrix such that we have a dressed state far from the initial state. Hence, once our protocol induces quicker relaxation towards equilibrium, these two features trigger genuine quantum Mpemba effect. In Sec. V, we illustrate our findings for a two-level system, and for the dynamics of the transverse field Ising model, and the XXZ model, each of these spin chains weakly coupled to a thermal bath. In Sec. VI, we showed that there are initial states for which an exponentially rapid relaxation to the equilibrium is already observed, but which are not incoherent with respect to the energy eigenbasis. Finally, in Sec. VII we summarize our conclusions.
@@ III Speeding up the relaxation process
The unitary transformation in Eq. (III) first maps $\rho({t_{0}})$ to its diagonal form. The permutation matrix $P_{\pi}$} then rearranges the order of the diagonal entries of $D$, which in turn is mapped onto the diagonal matrix ${P_{\pi}}{D}{P_{\pi}^{\dagger}}=\text{diag}({\lambda_{\pi(1)}},\ldots,{\lambda_{\pi(d)}})$. We note that $P_{\pi}$} is constructed in order to maximize the distinguishability between ${\rho^{\prime}}({t_{0}})$ and the steady state $\texttt{R}_{1}$. This means that the diagonal entries of $D$ can be rearranged in ascending order, ${\lambda_{\pi(1)}}<\ldots<{\lambda_{\pi(d)}}$, or even descending order, ${\lambda_{\pi(1)}}>\ldots>{\lambda_{\pi(d)}}$, depending on the figure of merit used to evaluate the distance from the transformed state to equilibrium. It will become clear later that $P_{\pi}$} plays an essencial role in ensuring that the genuine the Mpemba effect occurs for a given nonunitary dynamics [see Sec. IV]. Finally, the resulting matrix is rotated by $U_{1}$ to the quantum state in Eq. (III), which is incoherent with respect to the eigenbasis of the Hamiltonian $H$. In comparison with the results from Ref. Moroder et al. (2024), we note that the numerical simulation of our protocol requires low computational effort. The calculations involve prior knowledge of the spectrum of operators $H$ and $\rho({t_{0}})$, while the permutation matrix is determined by the appropriate ordering of the eigenvalues of $\rho({t_{0}})$ to maximize the distance from ${\rho^{\prime}}({t_{0}})$ to equilibrium. We note that, for higher dimensional systems, the spectral analysis of the Liouvillian can be implemented using the results presented in Ref. Westhoff et al. .
@@
We observe that the operators $\mathcal{L}_{\text{vec}}$} and $\Xi$ share the same spectrum of $d^{2}$} eigenvalues, of which $d$ are real and come from the subblock $\Xi_{\text{off}}$}, and the remaining d(d1)d(d-1) eigenvalues are complex and related to $\Xi_{\text{diag}}$}. On the one hand, we find that the subblock $\Xi_{\text{off}}$} contributes with $d$ eigenvectors $|{\texttt{L}^{\prime}_{s}}\rangle\rangle$ and $|{\texttt{R}^{\prime}_{s}}\rangle\rangle$, whose matrix forms are Hermitian, fully diagonal eigenmatrices $\texttt{L}^{\prime}_{s}$} and $\texttt{R}^{\prime}_{s}$} respective to the basis $\{|{k}\rangle\}_{k=1,\ldots,d}$. This naturally includes the eigenoperator $\texttt{R}^{\prime}_{1}$} related to the steady state ${\texttt{R}_{1}}={U_{1}}{\texttt{R}^{\prime}_{1}}{U_{1}^{\dagger}}$ of the dynamics, and also the vector $|{\texttt{L}^{\prime}_{1}}\rangle\rangle$ constrained to the eigenmatrix $\texttt{L}_{1}$} that plays the role of the identity. On the other hand, the slowest decaying modes of the nonunitary dynamics are related to the subspace spanned by d(d1)d(d-1) eigenvectors of $\Xi_{\text{diag}}$}. Noteworthy, the respective set of left $|{\texttt{L}^{\prime}_{s}}\rangle\rangle$ and right $|{\texttt{R}^{\prime}_{s}}\rangle\rangle$ eigenvectors gives rise to upper/lower triangular matrices $\texttt{L}^{\prime}_{k}$ and $\texttt{R}^{\prime}_{s}$}, respectively, which have all entries equal to zero except for one off-diagonal element. In detail, one finds that $\langle{j}|{\texttt{L}^{\prime}_{s}}|{l}\rangle={\delta_{j,{j_{0}}}}{\delta_{l,{l_{0}}}}$, and $\langle{j}|{\texttt{R}^{\prime}_{s}}|{l}\rangle={\delta_{j,{j_{0}}}}{\delta_{l,{l_{0}}}}$, for a given pair ${j_{0}},{l_{0}}\in\{1,\ldots,d\}$, with ${j_{0}}\neq{l_{0}}$. In particular, we note that this set comprises the eigenoperator ${\texttt{L}^{\prime}_{2}}={U_{1}^{\dagger}}{\texttt{L}_{2}}{U_{1}}$ connected to the lowest decaying mode of the Davies map. It turns out that ${\texttt{L}^{\prime}_{2}}=|{j_{0}}\rangle\langle{l_{0}}|$ is a non-Hermitian, upper/lower triangular matrix.
@@ IV.1 Hilbert-Schmidt distance
${\mathcal{D}_{\text{HSD}}}(\rho,\varrho)=\sqrt{\text{Tr}[{(\rho-\varrho)^{2}}]}\penalty 10000\ .$ (17)
We note that HSD is (i) non-negative, i.e., ${\mathcal{D}_{\text{HSD}}}(\rho,\varrho)\geq 0$, with ${\mathcal{D}_{\text{HSD}}}(\rho,\varrho)=0$ if and only if $\rho=\varrho$; (ii) symmetric, ${\mathcal{D}_{\text{HSD}}}(\rho,\varrho)={\mathcal{D}_{\text{HSD}}}(\varrho,\rho)$; (iii) isometric invariant, ${\mathcal{D}_{\text{HSD}}}(V\rho{V^{\dagger}},V\varrho{V^{\dagger}})={\mathcal{D}_{\text{HSD}}}(\rho,\varrho)$, where ${V^{\dagger}}={V^{-1}}$ is a unitary operator. It is worth to note that HSD satisfy the triangle inequality, ${\mathcal{D}_{\text{HSD}}}(\rho,\varrho)\leq{\mathcal{D}_{\text{HSD}}}(\rho,\varpi)+{\mathcal{D}_{\text{HSD}}}(\varpi,\varrho)$, for all $\rho,\varpi,\varrho\in\mathcal{S}$. However, HSD it is not always contractive under completely positive and trace-preserving maps Ozawa (2000).
@@
where $A_{l}$ is a permutation matrix, $0\leq{\epsilon_{l}}\leq 1$, with ${\sum_{l}}\,{\epsilon_{l}}=1$, while $D=\text{diag}({\lambda_{1}},\ldots,{\lambda_{d}})$ and $\Sigma=\text{diag}({\alpha_{1}},\ldots,{\alpha_{d}})$ are diagonal matrices containing the spectrum of $\rho({t_{0}})$ and $\texttt{R}_{1}$}, respectively. The right-hand side of Eq. (19) defines a convex sum ${\sum_{l}}\,{\epsilon_{l}}{c_{l}}$ of the nonnegative elements ${c_{l}}:=\text{Tr}({A_{l}}{D}{A_{l}^{\dagger}}\Sigma)$. We note that ${c_{\text{min}}}\leq{\sum_{l}}\,{\epsilon_{l}}{c_{l}}\leq{c_{\text{max}}}$, where ${c_{\text{min}}}=\text{min}\{{c_{1}},\ldots,{c_{d}}\}$ and ${c_{\text{max}}}=\text{max}\{{c_{1}},\ldots,{c_{d}}\}$. This means that there exist an optimal permutation matrix $P_{\pi,\text{opt}}$} such that $\text{Tr}({P_{\pi,\text{opt}}}{D}{P_{\pi,\text{opt}}^{\dagger}}\Sigma)$ is closest to the minimum value $c_{\text{min}}$, i.e., $\text{Tr}({P_{\pi,\text{opt}}}{D}{P_{\pi,\text{opt}}^{\dagger}}\Sigma)\leq{c_{\text{min}}}$. To see this point, we note that
@@
where $\pi(\bullet)$ maps index $l$ to $\pi(l)$, with $l,\pi(l)\in\{1,\ldots,d\}$. Therefore, to ensure that ${\sum_{l=1}^{d}}\,{\alpha_{l}}{\lambda_{\pi(l)}}\leq{c_{\text{min}}}$, it suffices to choose a permutation matrix such that ${P_{\pi,\text{opt}}}{D}{P^{\dagger}_{\pi,\text{opt}}}=\text{diag}({\lambda_{\pi(1)}},\ldots,{\lambda_{\pi(d)}})$, with ${\lambda_{\pi(1)}}\leq\ldots\leq{\lambda_{\pi(d)}}$ being listed in ascending order, provided that ${\alpha_{1}}\geq\ldots\geq{\alpha_{n}}$ are listed in descending order, or vice-versa Hardy et al. (1988). This proves the validity of the inequality in Eq. (18), which means that there exists a permutation $P_{\pi}$ that maximizes HSD evaluated for states ${\rho^{\prime}}({t_{0}})$ and $\texttt{R}_{1}$}.